# Protap: A Benchmark for Protein Modeling on Realistic Downstream Applications

### v2

@@ C.3 Protein–Ligand Interactions
Settings. For this task, we retain the same architectural hyperparameter configuration as used during pretraining. To encode SMILES representations, we adopt the GVP [10] architecture with a node dimensionality of 128, an edge dimensionality of 32, and a total of 3 layers. The maximum length of the pocket amino acid sequence is set to 85 residues, with shorter sequences padded and longer sequences truncated accordingly. Models are trained for 50 epochs with a learning rate of 5e-4, a batch size of 48 or 96, and the Adam optimizer. We report the mean and standard deviation of performance across a fixed set of random seeds. All experiments are conducted on NVIDIA L40 GPUs, with per-run training times ranging from approximately 20 minutes to 5 hours, depending on the model and task complexity.
@@
  • Original Data Source. The original data was collected from the KDBNet [43]. The data is stored at https://www.dropbox.com/s/owc45bzbfn05ix4/data.tar.gz. Based on the existing DAVIS [63] and KIBA [64] datasets, the authors further extracted the binding pockets of proteins from protein-ligand complexes for use in modeling this task.
@@
    • Following the approach in [62], we convert the format of the protein structural data provided by the authors and store it as a dictionary in a JSON file:
@@
                  "seq": "EVLAEGGFAIVFLCALKRMVCKREIQIMRDLS...",
                  "coord": [[[6.6065,16.2524,52.3289], ...], ...]
[…]
@@
    Table 14: Statistics of protein-ligand datasets.
@@
  • Hosting. A preprocessed version of the dataset can be downloaded from https://huggingface.co/datasets/findshuo/Protap/tree/main/PLI_DAVIS.
@@
C.4 Protein Function Annotation Prediction
@@
=$\mathcal{L}=\{\text{BP},\text{MF},\text{CC}\}$
@@
Settings. For this task, we adopt the same architectural hyperparameter settings as used during pretraining to ensure consistency. Each model is trained for 50 epochs using the Adam optimizer with a learning rate of 5e-4 and a batch size of 24. To ensure robustness, we report the average performance and standard deviation across a predefined set of random seeds. All training is conducted on NVIDIA L40 GPUs, with individual runs taking between 2 minutes and 10 hours, depending on the model architecture and task complexity.
@@
  • Data Source. The original data are from DeepFRI [45](https://github.com/flatironinstitute/DeepFRI), and the corresponding structural data are collected by ProteinWorkshop [27](https://zenodo.org/records/8282470/files/GeneOntology.tar.gz?download=1).
@@
          "seq": "SKKAELAELVKELAVYVDLRRATLHARASRLIGELLRELTADWDYVA...",
          "coords": [[[6.4359, 51.3870, 15.4490], ...], ...],
          "molecular_function": [0, 0, ...1,...],
          "biological_process": [0, 0, ...0,...],
          "cellular_component": [0, 0, ...1,...]
@@
    Table 15: Statistics of positive and negative functional labels in the training and test sets.
    Objective #Train #Test Label: 1 (Train) Label: 0 (Train) Label: 1 (Test) Label: 0 (Test)
    Number 23760 202 121,650 11,496,990 17,709 971,538
  • Hosting. A preprocessed version of the dataset can be downloaded from https://huggingface.co/datasets/findshuo/Protap/tree/main/AFP.
@@
C.5 Mutation Effect Prediction for Protein Optimization
Task definition. The log-ratio between wild-type and mutant amino acid probabilities has been shown to be an effective estimator of mutational impact [65, 66]. In the zero-shot setting, we do not access any label information. Instead, we perform inference using a model pretrained with masked language modeling (MLM). The goal is to quantify the log-likelihood of protein variants under the background. The calculation is shown in the equation.
@@ C.5 Mutation Effect Prediction for Protein Optimization
Settings. In the zero-shot setting, we do not use any label information, nor do we perform further training or fine-tuning. The prediction of mutation effects relies on the model’s estimated probabilities for the mutant and wild-type amino acids at the mutation site. Therefore, this setting is only applicable to models pre-trained using the masked language modeling (MLM) approach. Other types of pretraining methods are not suitable for this zero-shot scenario, and no additional domain-specific models are introduced in this application.
@@
  • Data Source. The dataset we used to evaluate in this benchmark is from ProteinGym [45](https://proteingym.org/)
  • Pre-process. Structural data is predicted by OmegaFold [67], where the input sequences for structure prediction are the wild-type sequences. We removed samples for which OmegaFold failed to generate structural predictions due to excessive sequence length.
@@
         F1I      ITLIELMIVIAIVGILAAVALPAYQDYTA...     -3.598         0
         F1L      LTLIELMIVIAIVGILAAVALPAYQDYTA...     -0.678         0
         F1V      VTLIELMIVIAIVGILAAVALPAYQDYTA...      1.299         1
         F1S      STLIELMIVIAIVGILAAVALPAYQDYTA...     -0.127         0
         T2A      FALIELMIVIAIVGILAAVALPAYQDYTA...      0.786         1
@@
    Table 16: Overall statistics of the ProteinGym DMS substitution dataset.
@@
Appendix D Domain-Specific Model Architecture
[…]
$h^{(\ell)}_{i}=W_{1}^{(\ell)}h^{(\ell-1)}_{i}+\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\left(W_{2}^{(\ell)}h^{(\ell-1)}_{j}+W_{3}^{(\ell)}e_{ij}\right).$ (1)
With attention weights defined as:
$\alpha_{ij}=\text{softmax}\left(\frac{\left(W_{4}^{(\ell)}h_{i}^{(\ell-1)}\right)^{\top}\left(W_{5}^{(\ell)}h_{j}^{(\ell-1)}+W_{3}^{(\ell)}e_{ij}\right)}{\sqrt{d_{\ell}}}\right).$ (2)
DeepPROTACs. DeepPROTAC [23] is a deep learning framework designed to predict the degradation efficacy of PROTAC molecules by integrating structural and chemical information from multiple components: the protein of interest (POI), the E3 ligase, and the PROTAC compound itself. The input includes five parts: the POI binding pocket, the E3 ligase binding pocket, the warhead (the PROTAC moiety binding to the POI), the E3 ligand (binding to the E3 ligase), and the linker represented as a SMILES string. For the four molecular structures (POI pocket, E3 pocket, warhead, and E3 ligand), atom-level graphs are constructed and processed using Graph Convolutional Layers (GCLs) followed by max pooling to obtain fixed-size feature embeddings. The linker is embedded using an LSTM network to capture sequential chemical features. All five embeddings are concatenated and passed through fully connected layers to yield a binary output: 1 indicates good degradation (defined as $\mathrm{DC}_{50}\leq 100\,\mathrm{nM}$ and $\mathrm{D}_{\max}\geq 80\%$), while 0 indicates poor degradation ($\mathrm{DC}_{50}>100\,\mathrm{nM}$ or $\mathrm{D}_{\max}<80\%$). This modular architecture enables DeepPROTACs to effectively model ternary complex formation and predict degradation outcomes based on both structural and sequential representations.
ET-PROTACs. ET-PROTACs [24] is an end-to-end deep learning model designed for PROTAC degradation prediction, consisting of four main components: (i) initial PROTAC and protein featurization; (ii) representation learning using 3D graph-based and sequence-based encoders; (iii) a cross-modal ternary attention block; and (iv) a final classifier. Each component is described in detail below. Each PROTAC is represented as a 2D molecular graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and associated with a 3D coordinate matrix $X\in\mathbb{R}^{|\mathcal{V}|\times 3}$ using RDKit [^3] https://www.rdkit.org/. Nodes $v_{i}\in\mathcal{V}$ represent atoms, and edges $a_{ij}\in\mathcal{E}$ denote chemical bonds. Node features $h_{i}$ and edge features $a_{ij}$ are encoded using learned embeddings. Additionally, atomic 3D coordinates $x_{i}\in\mathbb{R}^{3}$ are embedded to capture spatial information. The input protein is given as a sequence $P=\{a_{1},a_{2},\ldots,a_{n}\}$ of 23 amino acids (including one non-standard residue). Protein sequences are encoded by two embedding layers: a learned character embedding of dimension 64, and a positional embedding of the same dimension. The molecular graph and 3D coordinates are processed by an Equivariant Graph Neural Network (EGNN), which is invariant to rotations and translations. Let $h_{i}^{\ell}$, $x_{i}^{\ell}$, and $a_{ij}$ be the node features, coordinates, and edge features at layer $\ell$, respectively. The EGNN updates are as follows:
$m_{ij}$ $=\phi_{e}(h_{i}^{\ell},h_{j}^{\ell},\|x_{i}^{\ell}-x_{j}^{\ell}\|^{2},a_{ij}),$ (1)
$m_{i}$ $=\sum_{j\in\mathcal{N}(i)}m_{ij},$
$h_{i}^{\ell+1}$ $=\phi_{h}(h_{i}^{\ell},m_{i}),$
where $C=1/(M-1)$, $\phi_{e}$, $\phi_{x}$, and $\phi_{h}$ are learnable functions, and $M$ is the number of atoms. Final PROTAC embeddings are obtained by combining features of the warhead, linker, and E3 ligase substructures. ET-PROTACs leverages CNNs to encode protein sequences directly from FASTA format. The sequence embedding $E$ is passed through three CNN layers to yield latent protein features $P_{\text{cnn}}\in\mathbb{R}^{d\times f}$, where $f$ is the number of filters in the last layer. Given feature matrices $D=\{d_{1},\ldots,d_{X}\}$ (PROTAC), $P=\{p_{1},\ldots,p_{Y}\}$ (protein), and $L=\{l_{1},\ldots,l_{Z}\}$ (ligase), we form composite pairs:
$\text{pd}_{i}$ $=\text{CONCAT}(p_{j},d_{i}),$
$\text{dl}_{j}$ $=\text{CONCAT}(d_{i},l_{k}),$
$\text{pda}_{i}$ $=F(W_{\text{pd}}\cdot\text{pd}_{i}+b),$
$\text{dla}_{j}$ $=F(W_{\text{dl}}\cdot\text{dl}_{j}+b),$
$A_{i,j}$ $=F(W_{a}\cdot(\text{pda}_{i}+\text{dla}_{j})+b),$
$A_{\text{pd}}$ $=\sigma\left(\text{MEAN}(A,2)\right),$
$A_{\text{dl}}$ $=\sigma\left(\text{MEAN}(A,1)\right),$
[…]
DPFunc. DPFunc [21] is a structure-aware protein function prediction framework that captures residue-level information by integrating 3D geometric structure and pretrained sequence embeddings. For a given protein of length $l$, a residue-level undirected graph is constructed where each node represents a residue, and an edge is added between two residues if the distance between their Cα atoms is less than 10 Å. This defines the adjacency matrix $A\in\{0,1\}^{l\times l}$. Node features are initialized using embeddings from a pretrained protein language model (ESM-1b), resulting in $X\in\mathbb{R}^{l\times d}$. Two graph convolutional layers are then used to propagate structural information, with residual connections added to facilitate gradient flow. The update rule for the $k$-th GCN layer is given by:
$X^{(k+1)}=X^{(k)}+\text{ReLU}\left(\tilde{D}^{-1/2}\tilde{A}\tilde{D}^{-1/2}X^{(k)}W^{(k)}\right),$ (1)
$H=\text{ReLU}\left(\left(\text{ReLU}(IPR\cdot W_{\text{emd}})W_{1}+b_{1}\right)W_{2}+b_{2}\right).$ (2)
$Q_{i}=HW^{Q}_{i},\quad K_{i}=X_{\text{final}}W^{K}_{i},\quad V_{i}=X_{\text{final}}W^{V}_{i},$ (3)
and attention weights are derived as:
$W^{A}_{i}=\text{Softmax}\left(\frac{K_{i}Q_{i}^{\top}}{\sqrt{d}}\right).$ (4)
The multi-head attention output is computed as:
$X_{\text{multi}}=\text{LayerNorm}\left(\text{Concat}(W^{A}_{1}V_{1},\ldots,W^{A}_{n}V_{n})W_{\text{trans}}+X_{\text{final}}\right),$ (4)
followed by a feedforward transformation with residual connection:
$X_{\text{out}}=\text{LayerNorm}\left(\text{FF}(X_{\text{multi}})+X_{\text{multi}}\right).$ (5)
$x_{\text{pool}}=\sum_{i=1}^{l}X_{\text{out}}[i].$ (6)
$x_{\text{integrate}}=\text{Concat}\left(x_{\text{pool}},\frac{1}{l}\sum_{i=1}^{l}X[i]\right),$ (7)
$\hat{y}=\text{Sigmoid}(\text{MLP}(x_{\text{integrate}})).$ (8)
$\hat{y}^{\text{post}}_{i}=\max\left(\hat{y}_{i},\hat{y}_{\text{child}_{1}},\ldots,\hat{y}_{\text{child}_{n}}\right).$ (9)
UniZyme. UniZyme [22] is a biochemically informed framework designed to generalize protein cleavage site prediction across diverse enzymes. It comprises an enzyme encoder and a substrate encoder. The enzyme encoder integrates sequence-derived features with energetic frustration and 3D structural information. Given an enzyme $P_{e}=(X,R)$, where $X$ are residue embeddings and $R$ are Cα coordinates, the pairwise frustration score is defined as:
$F(i,j)=\frac{E(i,j)-\mu_{\text{rand}}(i,j)}{\sigma_{\text{rand}}(i,j)}.$ (1)
$\Phi^{\text{energy}}_{i,j}=\text{MLP}(\phi_{\text{energy}}(F(i,j))),\quad\Phi^{\text{dist}}_{i,j}=\text{MLP}(\phi_{\text{dist}}(\|r_{i}-r_{j}\|_{2})).$ (2)
$A^{k}_{i,j}=\frac{(h^{k-1}_{i}W_{Q})(h^{k-1}_{j}W_{K})^{T}}{\sqrt{d}}+\Phi^{\text{energy}}_{i,j}+\Phi^{\text{dist}}_{i,j}.$ (3)
$\hat{a}_{i}=\sigma(h_{i}\cdot w_{a}).$ (4)
$h_{e}=\sum_{i=1}^{N}\frac{f(\hat{a}_{i})}{\sum_{j}f(\hat{a}_{j})}h_{i}.$ (5)
$\hat{c}_{e,s}^{(t)}=\text{MLP}(\text{CONCAT}(H_{s}^{t:t+l},h_{e})).$ (6)
$\mathcal{L}=\mathcal{L}_{c}(D_{c})+\lambda\mathcal{L}_{a}(D_{a}).$ (7)
MONN. MONN [44] is composed of four interconnected modules: a graph convolutional module for molecular representation, a CNN module for residue-level protein representation, a pairwise interaction module for atom-residue interaction estimation, and an affinity prediction module for compound-protein binding affinity estimation.
$v_{i}^{0}=f(W_{\text{init}}v^{\text{init}}_{i}),$ (1)
$t_{i}^{l}=\sum_{v_{k}\in\mathcal{N}(v_{i})}f(W^{l}_{\text{gather}}[v_{k}^{l-1},e_{i,k}]),$ (2)
followed by feature updates:
$u_{i}^{l}=f(W^{l}_{\text{update}}[t_{i}^{l},v_{i}^{l-1}]).$ (3)
Protein sequences are encoded by mapping residues to BLOSUM62 [68] columns and processed through a 1D CNN to obtain residue embeddings $\{r_{j}\}$ in $\mathbb{R}^{h_{1}}$.
$P_{i,j}=\sigma(f(W_{\text{atom}}v_{i})\cdot f(W_{\text{residue}}r_{j})).$ (4)
$h_{v,i}$ $=f(W_{v}v_{i}),\quad h_{s}=f(W_{s}s),\quad h_{r,j}=f(W_{r}r_{j}),$ (5)
$h_{c}$ $=\sum_{i}\alpha_{v,i}h_{v,i},\quad h_{p}=\sum_{j}\alpha_{r,j}h_{r,j}.$ (6)
The binding affinity is predicted by a regression on the outer product between $[h_{c},h_{s}]$ and $h_{p}$:
$a=W_{\text{affinity}}f(\text{flatten}([h_{c},h_{s}]\otimes h_{p})),$ (7)
where $W_{\text{affinity}}\in\mathbb{R}^{1\times 2h_{2}^{2}}$.
CLIPZyme. CLIPZyme [13] formulates enzyme screening as a retrieval task, where a predefined set of enzymes is ranked based on their predicted ability to catalyze a given chemical reaction. Each reaction $R$ and enzyme $P$ is encoded into a $d$-dimensional vector $r,p\in\mathbb{R}^{d}$ using learned encoders $f_{\text{rxn}}$ and $f_{\text{p}}$, respectively. A scoring function $s(r,p)$ computes the cosine similarity between the two embeddings:
$s_{ij}=s(r_{i},p_{j})=\frac{r_{i}\cdot p_{j}}{\|r_{i}\|\|p_{j}\|}.$ (1)
$\mathcal{L}_{ij}=-\frac{1}{2N}\left(\log\frac{e^{s_{ij}/\tau}}{\sum_{k}e^{s_{ik}/\tau}}+\log\frac{e^{s_{ij}/\tau}}{\sum_{k}e^{s_{kj}/\tau}}\right),$ (2)
To represent a reaction, atom-mapped molecular graphs of the reactants $G_{x}$ and products $G_{y}$ are first encoded by a directed message passing neural network [69] (DMPNN) $f_{\text{mol}}$, yielding atom features $a_{i}$ and bond features $b_{ij}$:
$a_{i},b_{ij}=f_{\text{mol}}(G_{x},G_{y}).$ (3)
$v_{i}^{(TS)}$ $=v_{i}^{(x)}=v_{i}^{(y)},$ (3)
$e_{ij}^{(TS)}$ $=b_{ij}^{(x)}+b_{ij}^{(y)}.$ (4)
$a^{\prime}_{i},b^{\prime}_{ij}$ $=f_{TS}(G_{TS}),$ (4)
$r$ $=\sum_{i}a^{\prime}_{i}.$ (5)
Appendix E Pretrain Model Architecture and Experimental Setup
E.1 Model Architecture
@@
Equivariant Graph Neural Networks (EGNN). EGNN [11] is designed to process graphs where each node is associated with both feature embeddings and spatial coordinates. EGNN preserves equivariance under Euclidean transformations (translation and rotation) and node permutations, making it well-suited for modeling molecular and protein structures.
@@
$\textbf{m}_{ij}$ $=\phi_{e}\left(\textbf{h}_{i}^{l},\textbf{h}_{j}^{l},\|\textbf{x}_{i}-\textbf{x}_{j}\|^{2},a_{ij}\right),$ (1)
$\textbf{x}_{i}^{l+1}$ $=\textbf{x}_{i}^{l}+C\sum_{j\neq i}(\textbf{x}_{i}^{l}-\textbf{x}_{j}^{l})\cdot\phi_{x}(\textbf{m}_{ij}),$ (2)
$\textbf{m}_{i}$ $=\sum_{j\neq i}\textbf{m}_{ij},$ (3)
$\textbf{h}_{i}^{l+1}$ $=\phi_{h}(\textbf{h}_{i}^{l},\textbf{m}_{i}),$ (4)
@@
SE(3) Transformer. The SE(3) Transformer [41] is a neural architecture designed to model geometric data such as molecules and point clouds while respecting SE(3) symmetries, i.e., 3D rotations and translations. It achieves this through the integration of Tensor Field Networks (TFNs) [70] for equivariant message passing and invariant attention mechanisms for weighted aggregation. This section presents a unified formulation of the SE(3) Transformer, encompassing both TFN convolution and attention-based update steps.
Formally, each node $i$ in the graph is associated with a position $\textbf{x}_{i}\in\mathbb{R}^{3}$ and a set of typed features $\textbf{f}_{i}=\bigoplus_{\ell\geq 0}\textbf{f}_{i}^{\ell}$, where $\textbf{f}_{i}^{\ell}\in\mathbb{R}^{(2\ell+1)\times d}$ denotes a $-\ell$ feature tensor ($-\ell$ representation of SO(3)) with $d$ channels.
@@
$\textbf{f}^{\ell}_{\text{out},i}=\textbf{W}^{\ell\ell}_{\text{self}}\textbf{f}^{\ell}_{\text{in},i}+\sum_{k\geq 0}\sum_{j\in\mathcal{N}_{i}\setminus\{i\}}\alpha_{ij}\,\textbf{W}^{\ell k}(\textbf{x}_{j}-\textbf{x}_{i})\,\textbf{f}^{k}_{\text{in},j}.$ (1)
@@
$\textbf{W}^{\ell k}(\textbf{x})=\sum_{J=|\ell-k|}^{\ell+k}\varphi_{J}^{\ell k}(\|\textbf{x}\|)\sum_{m=-J}^{J}Y_{Jm}(\widehat{\textbf{x}})\,\textbf{Q}_{Jm}^{\ell k},$ (2)
where x^=x/$\|\textbf{x}\|$ = x / ∥ x ∥, $Y_{Jm}$ is the $m$-th spherical harmonic of order $J$, $\varphi_{J}^{\ell k}$ is a learnable radial function, and $\textbf{Q}_{Jm}^{\ell k}$ are learnable matrices formed from Clebsch–Gordan coefficients. This construction guarantees equivariance under SE(3) transformations by disentangling radial and angular dependencies.
@@
$\alpha_{ij}=\frac{\exp(\langle\textbf{q}_{i},\textbf{k}_{ij}\rangle)}{\sum_{j^{\prime}\in\mathcal{N}_{i}\setminus\{i\}}\exp(\langle\textbf{q}_{i},\textbf{k}_{ij^{\prime}}\rangle)},$ (3)
@@
$\textbf{q}_{i}=\bigoplus_{\ell\geq 0}\sum_{k\geq 0}\textbf{W}^{\ell k}_{Q}\textbf{f}^{k}_{\text{in},i},\quad\textbf{k}_{ij}=\bigoplus_{\ell\geq 0}\sum_{k\geq 0}\textbf{W}^{\ell k}_{K}(\textbf{x}_{j}-\textbf{x}_{i})\,\textbf{f}^{k}_{\text{in},j}.$ (4)
@@
[…]
@@
1:  $_{h}\leftarrow\textbf{W}_{h}\textbf{V}$ V
2:  $\textbf{V}_{\mu}\leftarrow\textbf{W}_{\mu}\textbf{V}_{h}$
3:  $\textbf{s}_{h}\leftarrow\|\textbf{V}_{h}\|_{2}$ (row-wise)
4:  $\textbf{v}_{\mu}\leftarrow\|\textbf{V}_{\mu}\|_{2}$ (row-wise)
5:  $_{h+n}\leftarrow\texttt{concat}(\textbf{s}_{h},\textbf{s})$ s )
6:  $\textbf{s}_{m}\leftarrow\textbf{W}_{m}\textbf{s}_{h+n}+\textbf{b}$
7:  $\textbf{s}^{\prime}\leftarrow\sigma(\textbf{s}_{m})$
8:  $\textbf{V}^{\prime}\leftarrow\sigma_{+}(\textbf{v}_{\mu})\odot\textbf{V}_{\mu}$
9:  return  $(\textbf{s}^{\prime},\textbf{V}^{\prime})$
@@
$h_{m}^{(j\to i)}$ $=\text{GVP}\left(\texttt{concat}(h_{v}^{(j)},h_{e}^{(j\to i)})\right),$ (1)
$h_{v}^{(i)}$ $\leftarrow\text{LayerNorm}\left(h_{v}^{(i)}+\frac{1}{k^{\prime}}\sum_{j\in\mathcal{N}_{i}}\text{Dropout}(h_{m}^{(j\to i)})\right),$ (2)

### v3

@@ F.3. Protein–Ligand Interactions
Settings. For this task, we retain the same architectural hyperparameter configuration as used during pretraining. To encode SMILES representations, we adopt the GVP (Jing et al., 2021) architecture with a node dimensionality of 128, an edge dimensionality of 32, and a total of 3 layers. The maximum length of the pocket amino acid sequence is set to 85 residues, with shorter sequences padded and longer sequences truncated accordingly. Models are trained for 50 epochs with a learning rate of 5e-4, a batch size of 48 or 96, and the Adam optimizer. We report the mean and standard deviation of performance across a fixed set of random seeds. All experiments are conducted on NVIDIA L40 GPUs, with per-run training times ranging from approximately 20 minutes to 5 hours, depending on the model and task complexity.
@@
  • Original Data Source. The original data was collected from the KDBNet (Luo et al., 2023b). The data is stored at https://www.dropbox.com/s/owc45bzbfn05ix4/data.tar.gz. Based on the existing DAVIS (Davis et al., 2011) and KIBA (Tang et al., 2014) datasets, the authors further extracted the binding pockets of proteins from protein-ligand complexes for use in modeling this task.
@@
    • Following the approach in (Ingraham et al., 2019), we convert the format of the protein structural data provided by the authors and store it as a dictionary in a JSON file:
@@
                  "seq": "EVLAEGGFAIVFLCALKRMVCKREIQIM..",
                  "coord": [[[6.60,16.25,52.32], ..], ..]
      For each structure, coords is a nested list of shape (N $\times$ 4 $\times$ 3), representing the 3D coordinates of the backbone atoms N, $\alpha, C$, and O for each residue in order, and $N$ is the length of the amino acid sequence.
@@
    Table 13. Statistics of protein–ligand datasets.
@@
F.4. Protein Function Annotation Prediction
@@
$\mathcal{L}=\{\text{BP},\text{MF},\text{CC}\}$
@@
Settings. For this task, to rigorously assess the model’s generalization ability, we included only those test sequences whose similarity to the training set was below 50%. we adopt the same architectural hyperparameter settings as used during pretraining to ensure consistency. Each model is trained for 50 epochs using the Adam optimizer with a learning rate of 5e-4 and a batch size of 24. To ensure robustness, we report the average performance and standard deviation across a predefined set of random seeds. All training is conducted on NVIDIA L40 GPUs, with individual runs taking between 2 minutes and 10 hours, depending on the model architecture and task complexity.
@@
  • Data Source. The original data are from DeepFRI (Gligorijević et al., 2021)(https://github.com/flatironinstitute/DeepFRI), and the corresponding structural data are collected by ProteinWorkshop (Jamasb et al., 2024)(https://zenodo.org/records/8282470/files/GeneOntology.tar.gz?download=1).
@@
          "seq": "SKKAELAELVKELAVYVDLRRATLHARASR...",
          "coords": [[[6.43, 51.38, 15.44], ..], ..],
          "molecular_function": [0, 0, ..1,..],
          "biological_process": [0, 0, ..0,..],
          "cellular_component": [0, 0, ..1,..]
@@
    Table 14. Statistics of positive and negative functional labels in the training and test sets.
    Objective #Train #Test Train Test
    Label: 1 Label: 0 Label: 1 Label: 0
    Number 23,760 202 121,650 11,496,990 17,709 971,538
@@
F.5. Mutation Effect Prediction for Protein Optimization
Task definition. The log-ratio between wild-type and mutant amino acid probabilities has been shown to be an effective estimator of mutational impact (Riesselman et al., 2018; Notin et al., 2022). In the zero-shot setting, we do not access any label information. Instead, we perform inference using a model pretrained with masked language modeling (MLM). The goal is to quantify the log-likelihood of protein variants under the background. The calculation is shown in the equation.
@@ F.5. Mutation Effect Prediction for Protein Optimization
Settings. In the zero-shot setting, we do not use any label information, nor do we perform further training or fine-tuning. The prediction of mutation effects relies on the model’s estimated probabilities for the mutant and wild-type amino acids at the mutation site. Therefore, pretraining methods whose models do not expose logits are unsuitable for this zero-shot setting, and no additional domain-specific models are incorporated in this study. Other types of pretraining methods are not suitable for this zero-shot scenario, and no additional domain-specific models are introduced in this application.
@@
  • Data Source. The dataset we used to evaluate in this benchmark is from ProteinGym (Gligorijević et al., 2021)(https://proteingym.org/)
  • Pre-process. Structural data is predicted by OmegaFold (Wu et al., 2022), where the input sequences for structure prediction are the wild-type sequences. We removed samples for which OmegaFold failed to generate structural predictions due to excessive sequence length.
@@
     F1I    ITLIELMIVIAIVGILA..   -3.598        0
     F1L    LTLIELMIVIAIVGILA..   -0.678        0
     F1V    VTLIELMIVIAIVGILA..   1.299         1
     F1S    STLIELMIVIAIVGILA..   -0.127        0
     T2A    FALIELMIVIAIVGILA..   0.786         1
@@
    Table 15. Overall statistics of the ProteinGym DMS substitution dataset.
@@
Appendix G Pretrain Model Architecture and Experimental Setup
G.1. Model Architecture
@@
Equivariant Graph Neural Networks (EGNN). EGNN (Satorras et al., 2021) is designed to process graphs where each node is associated with both feature embeddings and spatial coordinates. EGNN preserves equivariance under Euclidean transformations (translation and rotation) and node permutations, making it well-suited for modeling molecular and protein structures.
@@
(1) $\textbf{m}_{ij}$ $=\phi_{e}\left(\textbf{h}_{i}^{l},\textbf{h}_{j}^{l},\|\textbf{x}_{i}-\textbf{x}_{j}\|^{2},a_{ij}\right),$
(2) $\textbf{x}_{i}^{l+1}$ $=\textbf{x}_{i}^{l}+C\sum_{j\neq i}(\textbf{x}_{i}^{l}-\textbf{x}_{j}^{l})\cdot\phi_{x}(\textbf{m}_{ij}),$
(3) $\textbf{m}_{i}$ $=\sum_{j\neq i}\textbf{m}_{ij},$
(4) $\textbf{h}_{i}^{l+1}$ $=\phi_{h}(\textbf{h}_{i}^{l},\textbf{m}_{i}),$
@@
SE(3) Transformer. The SE(3) Transformer (Fuchs et al., 2020) is a neural architecture designed to model geometric data such as molecules and point clouds while respecting SE(3) symmetries, i.e., 3D rotations and translations. It achieves this through the integration of Tensor Field Networks (TFNs) (Thomas et al., 2018) for equivariant message passing and invariant attention mechanisms for weighted aggregation. This section presents a unified formulation of the SE(3) Transformer, encompassing both TFN convolution and attention-based update steps.
Formally, each node $i$ in the graph is associated with a position $\textbf{x}_{i}\in\mathbb{R}^{3}$ and a set of typed features $\textbf{f}_{i}=\bigoplus_{\ell\geq 0}\textbf{f}_{i}^{\ell}$, where $\textbf{f}_{i}^{\ell}\in\mathbb{R}^{(2\ell+1)\times d}$ denotes a rank-$\ell$ feature tensor (type-$\ell$ representation of SO(3)) with $d$ channels.
@@
(1) $\textbf{f}^{\ell}_{\text{out},i}=\textbf{W}^{\ell\ell}_{\text{self}}\textbf{f}^{\ell}_{\text{in},i}+\sum_{k\geq 0}\sum_{j\in\mathcal{N}_{i}\setminus\{i\}}\alpha_{ij}\,\textbf{W}^{\ell k}(\textbf{x}_{j}-\textbf{x}_{i})\,\textbf{f}^{k}_{\text{in},j}.$
@@
(2) $\textbf{W}^{\ell k}(\textbf{x})=\sum_{J=|\ell-k|}^{\ell+k}\varphi_{J}^{\ell k}(\|\textbf{x}\|)\sum_{m=-J}^{J}Y_{Jm}(\widehat{\textbf{x}})\,\textbf{Q}_{Jm}^{\ell k},$
where $\widehat{\textbf{x}}=\textbf{x}/\|\textbf{x}\|$, $Y_{Jm}$ is the $m$-th spherical harmonic of order $J$, $\varphi_{J}^{\ell k}$ is a learnable radial function, and $\textbf{Q}_{Jm}^{\ell k}$ are learnable matrices formed from Clebsch–Gordan coefficients. This construction guarantees equivariance under SE(3) transformations by disentangling radial and angular dependencies.
@@
(3) $\alpha_{ij}=\frac{\exp(\langle\textbf{q}_{i},\textbf{k}_{ij}\rangle)}{\sum_{j^{\prime}\in\mathcal{N}_{i}\setminus\{i\}}\exp(\langle\textbf{q}_{i},\textbf{k}_{ij^{\prime}}\rangle)},$
@@
(4) $\textbf{q}_{i}=\bigoplus_{\ell\geq 0}\sum_{k\geq 0}\textbf{W}^{\ell k}_{Q}\textbf{f}^{k}_{\text{in},i},\quad\textbf{k}_{ij}=\bigoplus_{\ell\geq 0}\sum_{k\geq 0}\textbf{W}^{\ell k}_{K}(\textbf{x}_{j}-\textbf{x}_{i})\,\textbf{f}^{k}_{\text{in},j}.$
@@
GVP. The Geometric Vector Perceptron (GVP) (Jing et al., 2021) is a neural module designed for learning over geometric data, in which each entity (e.g., an amino acid residue or atom) is represented by both scalar and vector features. Formally, given a pair $(\textbf{s},\textbf{V})$ where $\textbf{s}\in\mathbb{R}^{n}$ denotes scalar features and $\textbf{V}\in\mathbb{R}^{\nu\times 3}$ denotes geometric vector features, the GVP outputs a new pair $(\textbf{s}^{\prime},\textbf{V}^{\prime})\in\mathbb{R}^{m}\times\mathbb{R}^{\mu\times 3}$. This transformation is designed to ensure that the scalar outputs remain invariant while the vector outputs are equivariant with respect to rotations and reflections in $\mathbb{R}^{3}$.
@@
[…]
@@
(1) $h_{m}^{(j\to i)}$ $=\text{GVP}\left(\texttt{concat}(h_{v}^{(j)},h_{e}^{(j\to i)})\right),$
(2) $h_{v}^{(i)}$ $\leftarrow\text{LayerNorm}\left(h_{v}^{(i)}+\frac{1}{k^{\prime}}\sum_{j\in\mathcal{N}_{i}}\text{Dropout}(h_{m}^{(j\to i)})\right),$